### v2

# On novel Hamiltonian description of the nonholonomic Suslov problem

Abstract
We present some new Poisson bivectors that are invariants by the flow of the nonholonomic Suslov problem. Two rank four invariant Poisson bivectors have globally defined Casimir functions and, therefore, define cubic Poisson brackets on the five dimensional state space with standard symplectic leaves. For the Suslov gyrostat in the potential field we found rank two Poisson bivectors having only two globally defined Casimir functions and, therefore, we say about formal Hamiltonian description in these cases.
Dedicated to the memory of my friend and coathor Alexey Borisov
1 Introduction
The Suslov problem [23] describes the motion of a rigid body with a fixed point subject to a nonholonomic constraint that forces the angular velocity component along a given direction in the body to vanish. This nonholonomic system has been the subject of extensive research, see [2, 6, 9, 10, 11, 12, 13, 14, 15, 16, 21, 28] and references therein.
It allows us to start directly with a system of autonomous ordinary differential equations
$\dot{x}_{i}=X_{i}(x_{1},\ldots,x_{5})\,,\qquad i=1,\ldots,5\,$, (1.1)
where $x=(\gamma_{1},\gamma_{2},\gamma_{3},\omega_{1},\omega_{2})$ consists of three entries of the unit vector $\gamma$ and two entries of the angular velocity vector $\omega$, which are expressed in the special body frame. This frame is firmly attached to the body and its axes are arranged so that the nonholonomic constraint is
$\omega_{3}=0\,$,
whereas symmetric inertia tensor looks like
$I=\left(\begin{array}[]{ccc}I_{11}&0&I_{13}\\ 0&I_{22}&I_{23}\\ I_{13}&I_{23}&I_{33}\\ \end{array}\right)\,$,
with $I_{11}>0$ and $I_{22}>0$. One more condition must be fulfilled for real bodies [6].
The vector field $X$ is defined by the Euler-Poisson equations
$\dot{\gamma}_{1}=-\omega_{2}\gamma_{3},\qquad\dot{\gamma}_{2}=\omega_{1}\gamma_{3},\qquad\dot{\gamma}_{3}=\omega_{2}\gamma_{1}-\omega_{1}\gamma_{2}$
$I_{11}\dot{\omega}_{1}=-(I_{13}\omega_{1}+I_{23}\omega_{2})\omega_{2},\qquad I_{22}\dot{\omega}_{2}=(I_{13}\omega_{1}+I_{23}\omega_{2})\omega_{1}\,.$
See [6, 10] for a step-by-step derivation of these Euler-Poisson equations.
According to [10] solutions of the equations (LABEL:ep-eq) are meromorphic solutions iff either
$I_{13}=0,\qquad I_{11}=I_{22}+k^{2}\frac{I_{23}^{2}}{I_{22}}\,$,
or
$I_{23}=0,\qquad I_{22}=I_{11}+k^{2}\frac{I_{13}^{2}}{I_{11}}\,$,
where $k$ is a nonzero integer. In both cases exist three scalar solutions $f_{1}(x),f_{2}(x)$ and $f_{3}(x)$ of the invariance equation (1.5) which can be found in [10].
Below we study generic case without of these restrictions on entries of inertia tensor. Our aim is to discuss several previously unknown tensor invariants $T$ of the flow generated by $X$ (LABEL:ep-eq) which satisfy to the invariance equation
$\mathcal{L}_{X}\,T=0\,$, (1.3)
Here, $\mathcal{L}_{X}T$ is a Lie derivative of tensor field $T$ along vector field $X$ from (1.1) that determines the rate of change of the tensor field $T$ under the state space deformation defined by the flow of $X$. In local coordinates the Lie derivative of the tensor field $T$ of type $(p,q)$ is equal to
$({\mathcal{L}}_{X}T)^{i_{1}\ldots i_{p}}_{j_{1}\ldots j_{q}}=\sum_{k=1}^{n}X^{k}(\partial_{k}T^{i_{1}\ldots i_{p}}_{j_{1}\ldots j_{q}})$ $-\sum_{\ell=1}^{n}(\partial_{\ell}X^{i_{1}})T^{\ell i_{2}\ldots i_{p}}_{j_{1}\ldots j_{s}}-\ldots-\sum_{\ell=1}^{n}(\partial_{\ell}X^{i_{p}})T^{i_{1}\ldots i_{p-1}\ell}_{j_{1}\ldots j_{s}}$ (1.4)
$+\sum_{m=1}^{n}(\partial_{j_{1}}X^{m})T^{i_{1}\ldots i_{p}}_{mj_{2}\ldots j_{q}}+\ldots+\sum_{m=1}^{n}(\partial_{j_{q}}X^{m})T^{i_{1}\ldots i_{p}}_{j_{1}\ldots j_{q-1}m}$
where $\partial_{k}={\partial}/{\partial x_{k}}$ is the partial derivative on the $x_{k}$ coordinate.
The general theory of tensor invariants is discussed in the following classical books [22, 8] and in the modern review [18]. Different partial solutions of the invariance equation for integrable and non-integrable Hamiltonian systems are discussed in [24, 25, 26].
1.1 Known invariants
Let us describe known solutions of invariance equation (1.3) for the Suslov problem. In the space of scalar fields $f$ of type $(0,0)$ the invariance equation (1.3) has the form
$\mathcal{L}_{X}f=X^{1}\frac{\partial f}{\partial x_{1}}+\cdots+X^{5}\frac{\partial f}{\partial x_{5}}=0\,.$ (1.5)
We can solve this equation using the method of undetermined coefficients and obtain the energy and length of the Poisson vector $\gamma$ as a base in the space of solutions
$f_{1}=I_{11}\omega_{1}^{2}+I_{22}\omega_{2}^{2}\qquad\mbox{and}\qquad f_{2}=\gamma_{1}^{2}+\gamma_{2}^{2}+\gamma_{3}^{2}\,.$ (1.6)
Divergency of vector field $X$ is equal to
$\mbox{div}X=\sum_{k=1}^{5}\frac{\partial X_{k}}{\partial x_{k}}=\frac{I_{23}}{I_{22}}\omega_{1}-\frac{I_{13}}{I_{11}}\omega_{2}\,.$ (1.7)
Substituting divXX into the definition of the Darboux polynomial $D(x)$ as a cofactor
$\mathcal{L}_{X}D(x)=c(x)\cdot D(x)\,,\qquad c(x)=\mbox{div}X$ (1.8)
we obtain irreducible Darboux polynomial
$D(x)=I_{13}\omega_{1}+I_{23}\omega_{2}\,.$ (1.9)
If a solution of system (1.1) has a point on the hypersurface $D(x)=0$, then the whole solution is contained in this hypersurface, i.e. this invariant hypersurface divides the state space into invariant parts, which makes it easier to study the dynamics of the vector field $X$, see [20, 21].
Because cofactor $c(x)$ in (1.8) is equal to divergency we have an invariant measure
$\rho(x)=D(x)^{-1}$
which is singular on the invariant hypersurface $D(x)=0$, see discussion in [6, 10, 21]. So, there are invariant differential form of type (0,5)
$\Omega=D(x)^{-1}\,dx_{1}\wedge\cdots\wedge dx_{5}$ (1.10)
and invariant multivector field of type (5,0)
$\mathcal{E}=D(x)\,\frac{\partial}{\partial x_{1}}\wedge\cdots\wedge\frac{\partial}{\partial x_{5}}\,.$ (1.11)
Tensor products of invariants $\mathcal{E}df_{1,2}$ and $\mathcal{E}df_{1}df_{2}$ are also solutions of the invariance equation (1.3).
It is easy to see, that there is one more Darboux polynomial
$D_{im}(x)=\sqrt{-I_{11}\,}\omega_{1}-\sqrt{I_{22}\,}\omega_{2}$
so that
$\mathcal{L}_{X}D_{im}(x)=c_{im}(x)D_{im}(x)\,,\qquad c_{im}(x)=\sqrt{-\frac{1}{I_{11}I_{22}}\,}D(x)\,.$
The corresponding invariant hypersurface has no physical meaning.
2 Rank four invariant Poisson bivectors
The Poisson bivector $P$ is a contrvariant antisymmetric multivector fields of valency (2,0) which satisfy to the Jacoby identity
[[P,P]]=0[\![P,P]\!]=0 (2.1)
which we represent using the Schouten-Nijenhuis bracket [[.,.]][\![.,.]\!] on multivector fields. Any bivector field can be regarded as a skew homomorphism and the rank of this field at a point $x_{0}$ is the rank of the induced linear mapping [27].
Choosing local coordinates $x_{1},\ldots,x_{n}$ any Poisson bivector is given by
$P=\sum_{i<j}P^{ij}(x)\frac{\partial}{\partial x_{i}}\frac{\partial}{\partial x_{j}}\,$,
[…]
PdC(x)=0.PdC(x)=0\,.
The number of the independent Casimir functions of $P$ is related to the rank of $P$ [27].
A Poisson bracket is given in terms of $P$ by
$\{f,g\}=Pdfdg=\sum_{i<j}P^{ij}\frac{\partial f}{\partial x_{i}}\frac{\partial g}{\partial x_{j}},$
where $f$ and $g$ are functions on local coordinates $x$. In term of the Poisson bracket the Jacobi condition (2.1) looks like
$\{\{f,g\},h\}+\{\{g,h\},f\}+\{\{h,f\},g\}=0\,.$
The Poisson brackets between the Casimir function $C$ and any other function is equal to zero
$\{C,f\}=0\,,\qquad\forall f(x)\,.$
The Hamilton function $H(x)$ determines the Hamiltonian vector field $X$
$X=PdH\,,\qquad\mbox{or}\qquad X(g)=\{H,g\}\,.$ (2.2)
The Hamiltonian vector fields $X$ generate an integrable generalized distribution and the leaves of this foliation are symplectic. Usually the flow of the Hamiltonian vector field preserves the Poisson structure, it fixes each leaf and the Hamiltonian itself is a first integral.
So, we can try to compute Poisson bivector $P$ as a tensor invariant of a given vector field $X$, i.e. as a solution of invariance equation (1.3)
$\mathcal{L}_{X}P=0\,.$
The scalar invariant $H$ (2.2) is not fixed in this approach. It is computed later using the invariance property.
2.1 Cubic Poisson brackets
In the space of arbitrary multivector fields $T$ of valency $(2,0)$ the invariance equation looks like
$(\mathcal{L}_{X}T)^{ij}=\sum_{k=1}^{5}\left(X^{k}\frac{\partial T^{ij}}{\partial x^{k}}-T^{kj}\frac{\partial X^{i}}{\partial x^{k}}-T^{ik}\frac{\partial X^{j}}{\partial x^{k}}\right)=0\,.$ (2.3)
We solve this equation using the method of undetermined coefficients when we suppose that the entries of the multivector field $T(x)$ are inhomogeneous polynomials on variables $x_{1},\ldots,x_{5}$ of degree $N$.
At $N=1$ we have the following known rank two invariant bivector
$P_{1}=\left(\begin{array}[]{ccccc}0&\gamma_{3}&-\gamma_{2}&0&0\\ -\gamma_{3}&0&\gamma_{1}&0&0\\ \gamma_{2}&-\gamma_{1}&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{array}\right)\,.$ (2.4)
We use the notation $P_{1}$ instead of $T_{1}$ because this bivector satisfies the Jacobi condition (2.1), that allows as introduce the following linear Poisson bracket
$\{\gamma_{i},\gamma_{j}\}_{1}=\epsilon_{ijk}\gamma_{k}\,.$
Here $\epsilon$ is the totally antisymmetric Levi-Civita tensor.
At $N=2$ we have no solutions of the invariance equation (2.3).
At $N=3$ we have the following rank four invariant bivector
$T_{3}=(c_{1}f_{1}+c_{2}f_{2}+c_{3})P_{1}+c_{4}P_{2}+c_{5}P_{3}\,,\qquad c_{k}\in\mathbb{R}\,.$ (2.5)
Here $P_{1},P_{2}$ and $P_{3}$ are the Poisson bivectors which satisfy to the Jacobi condition (2.1). They are incompatible with each other and, therefore, $T_{3}$ is the symmetric skew multivector field of valence (2,0) which does not satisfy the Jacobi condition
$T_{3},T_{3}]\!]\neq 0\,.$
Nevertheless it is geometric structure which is invariant with respect to the Suslov model flow.
The entries of the Poisson bivectors $P_{2}$ and $P_{3}$ are equal to:
$P_{2}^{12}$ $=-\gamma_{3}(\omega_{1}\gamma_{1}+\omega_{2}\gamma_{2})\,,\qquad P_{2}^{13}=\gamma_{1}\gamma_{2}\omega_{1}-(\gamma_{1}^{2}+\gamma_{3}^{2})\omega_{2}\,,\qquad P_{2}^{23}=(\gamma_{2}^{2}+\gamma_{3}^{2})\omega_{1}-\gamma_{1}\gamma_{2}\omega_{2}\,$,
$P_{2}^{14}$ $=\phantom{-}I_{11}^{-1}\gamma_{1}\omega_{2}D(x)\,,\qquad P_{2}^{15}=-I_{22}^{-1}\gamma_{1}\omega_{1}D(x)\,,\qquad P_{2}^{24}=\phantom{-}I_{11}^{-1}\gamma_{2}\omega_{2}D(x)\,$,
$P_{2}^{25}$ $=-I_{22}^{-1}\gamma_{2}\omega_{1}D(x)\,,\qquad P_{2}^{34}=\phantom{-}I_{11}^{-1}\gamma_{3}\omega_{2}D(x)\,,\qquad P_{2}^{35}=-I_{22}^{-1}\gamma_{3}\omega_{1}D(x)\,$,
$P_{2}^{45}$ $=0\,$, (2.6)
and
$P_{3}^{12}$ $=0\,,\qquad P_{3}^{13}=0\,,\qquad P_{3}^{23}=0\,,\qquad P_{3}^{14}=-\omega_{1}\omega_{2}\gamma_{3}\,,\qquad P_{3}^{15}=-\omega_{2}^{2}\gamma_{3}\,$,
$P_{3}^{24}$ $=\omega_{1}^{2}\gamma_{3}\,,\qquad P_{3}^{2,5}=\omega_{1}\omega_{2}\gamma_{3}\,,\qquad P_{3}^{34}=\omega_{1}(\omega_{2}\gamma_{1}-\omega_{1}\gamma_{2})\,,\qquad P_{3}^{35}=\omega_{2}(\omega_{2}\gamma_{1}-\omega_{1}\gamma_{2})\,$,
$P_{3}^{45}$ $=-\frac{(I_{11}\omega_{1}^{2}+I_{22}\omega_{2}^{2})D(x)}{I_{11}I_{22}}\,.$ (2.7)
Here $D(x)$ is the Darboux polynomial (1.9).
Substituting generic solution $T_{3}$ (2.5) of the equation (2.3) into the Jacobi condition (2.1)
$T_{3},T_{3}]\!]=0$
and solving the resulting equations on coefficients $c_{1},\ldots,c_{5}$ we obtain the Poisson bivectors at
  1. 1.
    $c_{1}=0,\qquad c_{2}=c_{2},\qquad c_{3}=c_{3},\qquad c_{4}=0,\qquad c_{5}=c_{5}$;
  2. 2.
    $c_{1}=0,\qquad c_{2}=0,\qquad c_{3}=0,\qquad c_{4}=c_{4},\qquad c_{5}=c_{5}$;
  3. 3.
    $c_{1}=c_{1},\qquad c_{2}=c_{2},\qquad c_{3}=c_{3},\qquad c_{4}=0,\qquad c_{5}=0$;
  4. 4.
    $c_{1}=c_{1},\qquad c_{2}=0,\qquad c_{3}=0,\qquad c_{4}=2c_{5},\qquad c_{5}=c_{5}$.
In the first and last cases, the obtained Poisson bivectors have rank four. In the second and third cases, the corresponing Poisson bivectors have rank two.
In the first case the rank four invariant Poisson bivector is equal to
$P_{a}=(c_{2}f_{2}+c_{3})P_{1}+c_{5}P_{3}\,.$ (2.8)
A tensor operation between invariant tensor fields results in either a tensor invariant or zero. Therefore, we easy compute the following relations
$P_{a}df_{1}=2c_{5}f_{1}X\qquad\mbox{and}\qquad P_{a}df_{2}=0\,.$
So, function $f_{2}=|\gamma|^{2}$ is the globally defined Casimir function of $P_{a}$ and the original vector field $X$ has the standard Hamiltonian form
$X=P_{a}dH_{a}$
where Hamilton function
$H_{a}=\frac{1}{2c_{5}}\ln f_{1}\,$,
is the logarithm of energy up to a constant. If we set $c_{5}=1/4$, then we get $H_{a}=\ln f_{1}^{2}$. This is a globally defined function both for positive and negative values of energy.
In a similar way for the second rank four Poisson bivector
$P_{b}=c_{1}f_{1}P_{1}+2c_{5}P_{2}+c_{5}P_{3}$ (2.9)
we have
$P_{b}df_{1}=2c_{5}f_{1}X\,,\qquad P_{b}df_{2}=4c_{5}f_{2}X\,.$
It allows us to find the globally defined Casimir function
$C_{b}=\ln f_{1}^{2}+\ln f_{2}\,,\qquad P_{b}dC_{b}=0\,$,
since $f_{1}^{2}>0$ and $f_{2}>0$, and two equivalent Hamiltonian description of the original vector field
$X=P_{b}dH^{(1,2)}_{b}\,,\qquad H_{b}^{(1)}=\frac{\ln f_{1}}{2c_{5}}\,,\quad H_{b}^{(2)}=-\frac{\ln f_{2}}{4c_{5}}\,.$
3 Special case of the Euler-Poisson equations
The Euler equation describes the evolution of the angular velocity vector $\omega=(\omega_{1},\omega_{2},\omega_{3})$
$I\dot{\omega}+\omega\times I\omega=\mu r\times\gamma\,$,
Here $r=(r_{1},r_{2},r_{3})$ and $\gamma=(\gamma_{1},\gamma_{2},\gamma_{3})$ are radius vector of the center of mass and unit Poisson vector in the system of principal axes, $I=\mbox{diag}(I_{1},I_{2},I_{3})$ is a tensor of inertia, $\mu=mg$ is a weight of the body and $\times$ means a vector product, see all the details in textbook [3].
The Poisson equation describes evolution of the Poisson vector $\gamma$
$\dot{\gamma}=\gamma\times\omega\,$,
For a free rigid body $\mu=0$ and the Euler equation becomes independent of the Poisson equation. The Euler-Poisson equations for the Suslov problem (LABEL:ep-eq) are another example of equations that can be separated into two parts.
The third example is the Euler-Poisson equations
$I\dot{\omega}=\omega\times A\omega\qquad\mbox{and}\qquad\dot{\gamma}=\gamma\times B\omega\,$, (3.1)
where $A,B$ are two constant $3\times 3$ matrices. The system of equations (3.1) describes the partial motion of a rigid body with a fluid-filled cavity, see §3 of Chapter 3 in [3]. This is the case when the intensity of the fluid vortex in the cavity is small compared to the kinetic moment, or vice versa. These equations also arise in other physical applications, see [4, 5, 7] and references within.
The first equation in (3.1) is independent of the second equation which has zero divergency
$\sum_{i=1}^{3}\frac{\partial\left(\gamma\times BM\right)_{i}}{\partial\gamma_{i}}=0\,$,
so the phase flow of $X$ preserves the corresponding trivector $\mathcal{E}_{\gamma}$, vector field $Y$ and scalar invariant $F_{1}$ defined by
$\mathcal{E}_{\gamma}=\dfrac{\partial}{\partial\gamma_{1}}\wedge\dfrac{\partial}{\partial\gamma_{2}}\wedge\dfrac{\partial}{\partial\gamma_{3}}\,,\qquad Y=\sum_{i=1}^{3}\gamma_{i}\dfrac{\partial}{\partial\gamma_{i}}\,,\qquad F_{1}=(\gamma,\gamma)\,.$
If $A$ is a symmetric matrix, $A=A^{T}$, the divergency of the vector field $X$ (3.1) is equal to zero
$\mbox{div}\,X=\sum_{i=1}^{3}\frac{\partial\left(M\times AM\right)_{i}}{\partial M_{i}}+\sum_{i=1}^{3}\frac{\partial\left(\gamma\times BM\right)_{i}}{\partial\gamma_{i}}\,$,
and, therefore, the phase flow of $X$ preserves the volume form
$\Omega=d\gamma_{1}\wedge d\gamma_{2}\wedge d\gamma_{3}\wedge dM_{1}\wedge dM_{2}\wedge dM_{3}\,.$ (3.2)
In this case, the first equation in (3.1) coincides with the Euler equation describing free rigid body motion
$\dot{M}=M\times\omega\,,\qquad\mbox{where}\qquad\omega=AM\,$,
having two scalar invariants
$F_{2}=(M,M)\,,\qquad F_{3}=\frac{1}{2}(M,AM)\,.$
Substituting the well-known solutions $M(t)$ of the Euler equation into the second equation in (3.1) we non-autonomous Carathéodory differential equations of the form
$\dot{\gamma}=\gamma\times BM(t)\,.$
In the mathematical theory of control such systems are known as linear switched systems of differential equations.
According to the Euler–Jacobi–Lie theorem [17] if the system of ordinary differential equations
$\dot{x}=X(x_{1},\ldots,x_{n})$
on the $n$-dimensional smooth phase space with coordinates $x=(x_{1},\ldots,x_{n})$ possesses
  1. 1.
    $k$ functionally independent first integrals $F_{1},\ldots,F_{k}$;
  2. 2.
    m=nk2m=n-k-2 independent invariant symmetry fields $Y_{k+1},\ldots,Y_{n-2}$ which together with the field $Y_{n-1}=X$ generate a nilpotent Lie algebra of vector fields;
  3. 3.
    an invariant volume $n$-form $\Omega$;
    $\mathcal{L}_{X}\Omega=0\,.$
so that
$\mathcal{L}_{Y_{i}}F_{j}=0,\qquad\mathcal{L}_{Y_{i}}\Omega=0\,$,
then it is integrated by quadratures.
In our case, we have $k=3$ independent first integrals $F_{1}$, $F_{2}$ and $F_{3}$, m=632=1m=6-3-2=1 vector field $Y$ commuting with $X$ and invariant volume form $\Omega$ (3.2). However, the vector field $X$ (3.1) does not satisfy the conditions of the Euler–Jacobi–Lie theorem since
$\mathcal{L}_{Y}F_{1}=2F_{1}\,.$
So, in the generic case, we have a non-integrable system.
3.1 Formal and informal Hamiltonization
We have the following tensor invariants of $X$ (3.1) :
  • three scalar invariants
    $F_{1}=(\gamma,\gamma)\,,\qquad F_{2}=(M,M)\,,\qquad F_{3}=\frac{1}{2}(M,AM)\,;$ (3.3)
  • three invariant multivectors
    $Y=\sum_{i=1}^{3}\gamma_{i}\dfrac{\partial}{\partial\gamma_{i}}\,,\qquad\mathcal{E}_{\gamma}=\dfrac{\partial}{\partial\gamma_{1}}\wedge\dfrac{\partial}{\partial\gamma_{2}}\wedge\dfrac{\partial}{\partial\gamma_{3}}\,$, (3.4)
    and
    $\mathcal{E}=\dfrac{\partial}{\partial\gamma_{1}}\wedge\dfrac{\partial}{\partial\gamma_{2}}\wedge\dfrac{\partial}{\partial\gamma_{3}}\wedge\dfrac{\partial}{\partial M_{1}}\wedge\dfrac{\partial}{\partial M_{2}}\wedge\dfrac{\partial}{\partial M_{3}}\,.$
The Lie derivatives $\mathcal{L}_{X}$ of all these invariants are equal to zero. We will use these invariants to construct Hamilton representation of $X$
$X=PdH$
where $H$ is a Hamilton function and $P$ is a Poisson bivector.
Another main ingredient is the semi invariant or Darboux invariant $Z$ of the homogeneous vector field $X$
$\mathcal{L}_{X}Z=X\,$,
where vector field
$Z=\sum_{i=1}^{3}\gamma_{i}\dfrac{\partial}{\partial\gamma_{i}}+\sum_{i=1}^{3}M_{i}\dfrac{\partial}{\partial M_{i}}$ (3.5)
is the well-known Euler vector field whose flow is the linear multiplication action. Since the first integrals $F_{k}$ are homogeneous functions we have
$\mathcal{L}_{Z}F_{k}=2F_{k}\,,\qquad k=1,2,3,$
according to the Euler theorem on homogeneous functions. Moreover, the invariant rank-two bivector
$P_{e}=X\wedge Z\,,\qquad\mathcal{L}_{X}P=0\,$,
satisfies to the Jacobi condition
$P_{e},P_{e}]\!]=0\,$,
according to the same Euler theorem.
The invariant bivector $P$ does not have four independent Casimir functions since we have only three single-valued scalar invariants $F_{1},F_{2}$ and $F_{3}$, see discussion in [1]. So, we can say that this rank-two bivector yields formal Hamiltonization of $X$.
It is simple to confirm that the rank four bivector
$P=(c_{1}F_{1}+c_{2}F_{2}+c_{3}F_{3})P_{\gamma}+c_{4}P_{c}+eP_{e}\,,\qquad c_{k},e\in\mathbb{R}\,$, (3.6)
where
$P_{\gamma}=\mathcal{E}_{\gamma}dF_{1}\,,\qquad P_{c}=X\wedge Y\qquad\mbox{and}\qquad P_{e}=X\wedge Z\,.$
is the invariant bivector
$\mathcal{L}_{X}P=0\,$,
which satisfies the Jacobi condition
[[P,P]]=0,[\![P,P]\!]=0\,,
and the following equations hold
$PdF_{1}=2(c_{4}+e)F_{1}X\,,\qquad PdF_{2}=2eF_{2}X\,,\qquad PdF_{3}=2eF_{3}X\,.$ (3.7)
From (3.7) follows that we can define two independent Casimir functions
$C_{1}=\left(1+\frac{c_{4}}{e}\right)\ln F_{2}-\ln F_{1}\,,\qquad C_{2}=\left(1+\frac{c_{4}}{e}\right)\ln F_{3}-\ln F_{1}\,.$
This means that $P_{D}C_{1},2=0$, and the vector field $X$ (3.1) can be rewritten in Hamiltonian form
$X=PdH\,,\qquad H=\frac{1}{e}\ln F_{2}\,.$
Here $P$ is a rank four Poisson bivector with two well-defined Casimir functions $C_{1,2}$ and $F_{2}$ is a mechanical energy. We can therefore talk about the informal Hamiltonisation of the vector field $X$.
If $A$ is an arbitrary matrix $A\neq A^{T}$, the divergence of the vector field $X$ (3.1) does not equal zero.
$\mbox{div}\,X=(a_{2,3}-a_{3,2})M_{1}+(a_{3,1}-a_{1,3})M_{2}+(a_{1,2}-a_{2,})M_{3}$
and, therefore, the phase flow of $X$ does not preserve the volume form $\Omega$ (3.2) and function
$F_{3}=\frac{1}{2}(M,AM)\,.$
The rank four invariant Poisson bivector (3.6) now depends on four parameters $c_{k}$
$P=(c_{1}F_{1}+c_{2}F_{2})P_{\gamma}+c_{3}P_{c}+eP_{e}\,,\qquad c_{k},e\in\mathbb{R}\,.$ (3.8)
But we cannot now construct two independent single-valued Casimir functions $C_{1,2}$ and the Hamiltonian $H$ from only two scalar invariants $F_{1,2}$ of the phase flow. So, in this case, we can only say about the formal Hamiltonization of the vector field $X$.
4 Conclusion
Our aim is to consider a system of ordinary differential equations
$\dot{\mathrm{x}}_{k}=X_{k}(\mathrm{x}_{1},\ldots,\mathrm{x}_{N})\,.$
having the same structure as the Euler-Poisson equations for the free top, for the Suslov problem and for some other known systems of equations from the textbook [3]. For all these systems we can divide $N$ variables $\mathrm{x}=(\mathrm{x}_{1},\ldots,\mathrm{x}_{N})$ on two parts
$x=(x_{1},\ldots,x_{n})\qquad\mbox{and}\qquad y=(y_{1},\ldots,y_{m})\,,\qquad n+m=N$
and consider a system of equations with quadratic right-hand sides
$\dot{x}_{k}=\sum_{i,j=1}^{n}a_{k}^{ij}x_{i}x_{j}\qquad\mbox{and}\qquad\dot{y}_{\ell}=\sum_{i=1}^{n}\sum_{j=1}^{m}b_{\ell}^{ij}x_{i}y_{j}\,,\qquad a_{k}^{ij},b_{\ell}^{ij}\in\mathbb{R}\,.$ (4.9)
We also suppose that first subsystem of equations has n1n-1 independent first integrals $f_{1},\ldots,f_{n-1}$ depending on $x_{k}$, whereas second subsystem of equations has m2m-2 independent first integrals $g_{1},\ldots,g_{n-2}$ depending on $y_{\ell}$.
The intersection of the invariant levels of the first integrals $f_{1},\ldots,f_{n}$ determine solutions $x_{k}(t)$. When we substitute these solutions into the second subsystem, we obtain non-autonomous system of $m$ equations
$\dot{y}_{\ell}=\sum_{i=1}^{n}\sum_{j=1}^{m}b_{\ell}^{ij}x_{i}(t)y_{j}\,,\qquad\ell=1,\ldots,m,$
which has m2m-2 first integrals $g_{1},\ldots,g_{m-1}$. It allows us to study solutions of this counterpart of the Poisson equations in the framework of the Carathéodory theory.
We want to find an invariant Hamiltonian representation of differential equations (4.9). In this note we consider two partial cases of equations (4.9) with $N=5$ and $N=6$. A general case will be considered in subsequent publication.
Funding
The study was carried out with the financial support of the Ministry of Science and Higher Education of the Russian Federation in the framework of a scientific project under agreement No 075-15-2025-013 by St. Petersburg State University as part of the national project “Science and Universities” in 2025.